OpenStudy (anonymous):

Slove, writing any non-real solution in the form a+bi:x^4+10x^2=6x^3. Can you show details on how to answer please thank you for your help.

OpenStudy (campbell_st):

rewrite your equation \[x^4 - 6x^3 + 10x^2 = 0\] then factor \[x^2(x^2 -6x + 10) = 0\] so the non-real solutions occur in the quadratic factor. you need to solve the quadratic using the general quadratic formula and you'll find that the solutions will be in the form of x =a + bi hope it helps

OpenStudy (anonymous):

I have an answer key that shows the answer is 0,3+i,3-i. Can you show me how they got this answer please, thanks

OpenStudy (campbell_st):

well zero is correct and comes from the factor x^2 so looking at the quadratic you need \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] where a = 1, b = -6 and c = 10 substitute and evaluate then you'll see if you get the required answers...